| Operation Summary for March 2, 2012 - March 5, 2012 |
|
NuMI reported receiving 9.39E18 protons on target for the period from 2/27/12 to 3/5/12. Friday March 2 The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI and MiniBooNE experiments taking beam. During the day shift, Operations received the Run Conditions for the Neutrino Area E906 experiment (SeaQuest) at 8:51 AM, but Operations still needs the Operations Readiness Clearance (ORC) from PPD Safety before they can send the experiment beam. Switchyard (SY) personnel began an extraction study at 9:30 AM. Operators prepared the Neutrino Area enclosure N01/NM1 for access at 9:33 AM. At 9:46 AM, a Booster RF station (BRF5) tripped off repeatedly on a RF gap fault indication; a Booster RF person managed to resolve the problem. At 10:13 AM, a Linac Klystron RF station (KRF5) tripped off; it reset. Operators searched and secured the Meson MTest enclosures at 11:01 AM. At 11:17 AM, a Linac quadrupole magnet power supply tripped off; it reset. At 11:47 AM, Operations established beam to the Meson FTBF experiment T-1015. SY extraction studies ended at 11:47 AM. A Booster bulk power supply (B:CEP400) tripped off at 3:01 PM; it reset. During the evening shift, a MI study ended at 6:12 PM. At 7:13 PM, a Linac RF station (LRF3) tripped off for the first of two times in one minute; it reset. Operations disabled the T-1015 beam at 7:55 PM. At 10:17 PM, BRF9 and 10 tripped off on a vacuum indication for the first of two times in 12 minutes; they reset. Saturday March 3 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, at 3:06 AM, QPS108 tripped off; it reset. During the day shift, at 9:51 AM, LRF3 tripped off for the first of four times in two hours and 45 minutes; it reset. Operations established beam to T-1015 around 10:35 AM. At 11:04 AM, the MiniBooNE beam permit tripped off due to a MI-12 enclosure airflow alarm; it reset. At 1:59 PM, the MI-40 sump pump came into alarm; Operators prepared to access the enclosure, but the alarm cleared. During the evening shift, at 4:53 PM, the MI-40 sump pump alarmed again; Operators waited to see if the alarm would clear. It didn’t. Operators prepared MI-40 for access at 6:07 PM, accessed and found that the sump pit was starting to overflow; FESS personnel repaired the sump pump’s floats, but installed a backup pump in the pit because the floats were old and rusty and might fail again. Operations reestablished beam to MiniBooNE at 7:17 PM. Operators were trying to turn on MI after the accesses, but a MI quadrupole magnet power supply (I:Q703) was being held off on a volts to ground indication. Operators called in Power Supply personnel who found the load transfer switch was in the wrong mode. Unfortunately, there was still a problem the Q703 power supply that held off beam. Power Supply personnel replaced the cable driver card for Q703 and MI resumed sending beam around 10 PM. Operations disabled T-1015 beam at 10:02 PM. Sunday March 4 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, a P1 transfer line power supply (I:CPSF1) tripped off; it reset. During the day shift, at 10:47 AM, KRF1 tripped off; it reset. At 11:59 AM, B:CEP400 tripped off; it reset. Three MI RF ion pumps tripped off at 12:46 PM; they reset. During the evening shift, KRF3 tripped off at 4:18 PM; it reset. At 6:09 PM, a MI RF ion pump (I:IPRF22) tripped off and brought down BRF11 and 12; everything recovered and reset. At 6:54 PM, the NuMI beam permit tripped off due to repeated trips by a MI RF station (MIRF16); everything reset. LFR3 tripped off at 7:59 PM; it reset. At 8:34 PM, the MI South Anode Power Supply tripped off on an interlock fault indication; it reset. BRF16 tripped off at 10:23 PM; it reset. At 10:25 PM, the MiniBooNE beam permit tripped off twice due to an airflow alarm; a MiniBooNE person advised Operators on how to raise the alarm limit. At 11:33 PM, a SY Lambertson (S:MLAM2) tripped off on a magnet temperature indication and would not reset; Operators prepared SY enclosures C, D, and E for access to investigate and found the magnet cool. Power Supply personnel will repair it in the morning. Monday March 5 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, at 12:34 AM, QPS108 tripped off for the first of two times in about seven hours; it reset. At 4:38 AM, a LRF1 tripped off; it reset. A MI Kicker (I:KPS4A) tripped off on a low Fluorinert level; it reset. Operations Coordinator (OpsCo) The OpsCo reports that they will be a maintenance day on Wednesday, 3/7/12, to allow Booster RF personnel to upgrade BRF2. Beam will go off at 6 AM with keys available for accessing at 7 AM. |
|
More Information Comments and Suggestions |
| last modified 03/05/2012 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |